# Can Watermarked LLMs be Identified by Users via Crafted Prompts?

### v1

@@ Abstract
Text watermarking for Large Language Models (LLMs) has made significant progress in detecting LLM outputs and preventing misuse. Current watermarking techniques offer high detectability, minimal impact on text quality, and robustness to text editing. However, current researches lack investigation into the imperceptibility of watermarking techniques in LLM services. This is crucial as LLM providers may not want to disclose the presence of watermarks in real-world scenarios, as it could reduce user willingness to use the service and make watermarks more vulnerable to attacks. This work is the first to investigate the imperceptibility of watermarked LLMs. We design an identification algorithm called Water-Probe that detects watermarks through well-designed prompts to the LLM. Our key motivation is that current watermarked LLMs expose consistent biases under the same watermark key, resulting in similar differences across prompts under different watermark keys. Experiments show that almost all mainstream watermarking algorithms are easily identified with our well-designed prompts, while Water-Probe demonstrates a minimal false positive rate for non-watermarked LLMs. Finally, we propose that the key to enhancing the imperceptibility of watermarked LLMs is to increase the randomness of watermark key selection. Based on this, we introduce the Water-Bag strategy, which significantly improves watermark imperceptibility by merging multiple watermark keys.
@@ 1 Introduction
Text watermarking techniques for LLMs have become an important method to mitigate the above issues by adding detectable features to LLM outputs (Liu et al., 2024c). Recent researches on LLM watermarking have focused on improving watermark detectability (Kirchenbauer et al., 2023a), minimizing impact on generated text (Aaronson & Kirchner, 2022), and enhancing robustness against text modifications (Liu et al., 2024b). However, no work has considered the imperceptibility of watermarked LLMs, i.e., whether users can know if an LLM service is watermarked. In real-world scenarios, LLM service providers may not disclose the existence of watermarks, as it could reduce user willingness to use the service and make it more vulnerable to attacks (Sadasivan et al., 2023). As more LLM services consider implementing watermarks, it is crucial to investigate whether users can identify watermarked LLMs solely through crafted prompts.
Some studies focus on the imperceptibility of watermarked text, ensuring watermarked and non-watermarked texts are indistinguishable (Hu et al., 2023; Wu et al., 2023b). However, even if individual watermarked texts are imperceptible, the distribution of numerous watermarked texts may reveal whether the LLM is watermarked, especially when repeatedly sampling with the same watermark key (Wu et al., 2024). While some studies explore cracking watermarks using large volumes of watermarked text (Jovanović et al., 2024; Sadasivan et al., 2023; Wu & Chandrasekaran, 2024; Zhang et al., 2024b), they assume the target LLM is watermarked and cannot determine if the LLM is watermarked. Research on the imperceptibility of watermarked LLMs should investigate whether watermarked and non-watermarked LLMs can be distinguished.
@@
In this work, we conduct the first study on the imperceptibility of watermarked LLMs. We propose the Water-Probe algorithm to identify watermarked LLMs under two main types of watermarking methods. Our motivation stems from a key observation: current LLM watermarking algorithms expose consistent bias when repeatedly sampled under the same watermark key. Based on this, our Water-Probe algorithm first crafts prompts to perform repeated sampling under the same watermark key, then compares the consistency of sampling distribution differences across different prompts under a pair of watermark keys. Highly consistent differences indicate a watermarked LLM.
@@
Finally, we explore methods to enhance the imperceptibility of watermarked LLMs. We find that increasing the randomness of watermark key selection is crucial, as it makes it more difficult to construct prompts for repeated sampling using the same key. Based on this, we propose the Water-Bag algorithm, which combines multiple watermark keys into one, randomly selecting a key for each generation and choosing the highest score for detection. However, increasing key selection randomness often leads to a slight decrease in detectability. Addressing this trade-off should be an important direction for future work.
Our main contributions can be summarized as follows:
  • We conduct the first study on the imperceptibility of watermarked LLMs, proposing the Water-Probe algorithm to effectively identify watermarked LLMs by exploiting consistent biases in watermarking keys.
  • We demonstrate the high accuracy and robustness of our algorithm across various LLMs, watermarking methods, and generation settings, while maintaining a minimal false positive rate for non-watermarked LLMs.
  • We introduce the Water-Bag algorithm to enhance LLM watermark imperceptibility.
2 Related Work
Large language model (LLM) watermarking techniques (Liu et al., 2024c) have become crucial for copyright protection (Sander et al., 2024), generated text detection (Wu et al., 2023a), and preventing misuse (Liu et al., 2024d). These watermarking algorithms introduce subtle features during text generation that are almost imperceptible to humans but detectable by designed detection algorithms. LLM watermarking can be broadly categorized into two types. The first type is post-processing watermarking, which modifies generated text using format-based (Sato et al., 2023), lexical-based (Yang et al., 2023), syntax-based (Wei et al., 2022), or generation-based (Zhang et al., 2024a) approaches to add watermark. However, post-processing watermarking methods require waiting for text generation to complete before modifying and adding watermarks, which is not suitable for current LLM services that require real-time text generation. Another category of watermarking algorithms, known as watermarking during generation, typically involves adjusting the distribution of the next generated token based on a watermark key. For instance, the KGW (Kirchenbauer et al., 2023a) algorithm divides the vocabulary into red and green lists, increasing the probability of tokens in the green list. SIR (Liu et al., 2024b) further modifies logits based on semantic information, enhancing watermark robustness. UPV (Liu et al., 2024a) utilizes two independent networks for watermark embedding and detection, enabling public verification. Additionally, some research has explored watermarking techniques for code generation scenarios (Lu et al., 2024) or perform watermark detection on longer documents (Pan et al., 2024b) or even translated documents (He et al., 2024). An important objective of these methods is to maintain the imperceptibility of generated text, i.e., watermarked and non-watermarked text should have identical distributions, with some distortion-free algorithms showing promising results (Kuditipudi et al., 2023; Aaronson & Kirchner, 2022). However, previous work has overlooked the imperceptibility of watermarked LLMs themselves, i.e., whether external users can detect if an LLM service contains watermarks without disclosure. This work investigates the imperceptibility of watermarked LLMs.
@@ 3 Preliminaries
To facilitate subsequent discussion, we introduce LLMs and LLM watermarking algorithms, then present the objectives of this work. We first provide the definition of LLM used in this work as follows:
@@
N-gram based watermarking ensures that for the same n-token prefix, the watermark key for generating the next token remains consistent. This approach is widely used in current watermarking algorithms, including KGW (Kirchenbauer et al., 2023a), KGW-V2 (Kirchenbauer et al., 2023b), Aar (Aaronson & Kirchner, 2022), DiPmark (Wu et al., 2023b), and SIR (Liu et al., 2024b). Next, we define fixed-key-list-based watermarking:
@@ 4 Watermarked LLM Identification
In this section, we introduce our Water-Probe algorithm, which identifies watermarks in LLMs through well-designed prompts. We first analyze why watermarked LLM identification is possible, then present the general algorithm process as shown in Fig. 1, and finally describe two specific algorithm variants.
@@ 4.2 Pipeline of Watermarked LLM Identification
Observation 1 implies that the key to identifying a watermarked LLM is to construct prompts that allow for multiple samplings using the same watermark key to reveal the difference between $P_{M}^{F}(y_{i}|x,y_{1:i-1},k)$ and $P_{M}(y_{i}|x,y_{1:i-1})$. However, due to the black-box setting, we cannot directly access the origin logits $P_{M}(y_{i}|x,y_{1:i-1})$. Instead, we calculate the difference in LLM outputs for two distinct keys, defined as $\Delta(p,k_{m},k_{n})=P_{M}^{F}(y_{1}|p,k_{m})-P_{M}^{F}(y_{1}|p,k_{n})$, to aid in identification.
If two prompts yield similar output distributions, the same watermark key should have similar effects on both prompts (proven in Theorem 1). We determine if an LLM contains a watermark by comparing the consistency of the effects of two watermark keys on two similar prompts. Specifically, given $p_{1}$, $p_{2}$, $k_{1}$, and $k_{2}$, we assess the similarity between $\Delta(p_{1},k_{1},k_{2})$ and $\Delta(p_{2},k_{1},k_{2})$. High similarity indicates the presence of a watermark; otherwise, we conclude there is no watermark. Based on the above analysis, we now present the process of the Water-Probe algorithm.
Step 1: Construct highly correlated prompts. Construct $N$ prompts $P_{1},P_{2},...,P_{N}$ such that their output probability distributions under $M$ are highly similar, which can be expressed as:
$\forall i,j\in\{1,2,...,N\},\text{KL}(P_{M}(\cdot|P_{i})||P_{M}(\cdot|P_{j}))\leq\epsilon\text{ and }P_{i}\neq P_{j}$ (5)
@@
Step 2: Sampling with simulated fixed watermark keys. Since we cannot access the logits under a given watermark key, we need to use repeated sampling to estimate the distribution. We construct a set of simulated watermark keys $K=\{k_{1},k_{2},...,k_{m}\}$ based on our prompt design (detailed in subsequent sections). For each prompt $P_{i}$ and each simulated key $k_{j}\in K$, we estimate the distribution as follows:
$\hat{P}_{M}^{F}(y|P_{i},k_{j})=\frac{1}{W}\sum_{w=1}^{W}\mathbf{1}_{y_{i,j}^{w}=y},\quad\text{where }y_{i,j}^{w}\sim P_{M}^{F}(y|P_{i},k_{j})$ (6)
where $W$ is the sample count, $\mathbf{1}_{A}$ is the indicator function, and $y_{i,j}^{w}$ is the $w$-th sample sampled from $P_{M}^{F}(y|P_{i},k_{j})$. Specific prompt techniques for different watermarking algorithms will be detailed later. Note that our prompt design for simulating watermark keys assumes the target LLM has a watermark. If it doesn’t, then $P_{M}^{F}(y|P_{i},k_{j})=P_{M}(y|P_{i})$ for all simulated keys.
@@
For prompts $p_{1}$ and $p_{2}$ satisfying the similarity condition in Equation 5, the watermark rule $F$ satisfies Lipschitz continuity. That is, there exists a constant $L>0$ such that for the probability distributions $P_{M}(\cdot|p_{1}),P_{M}(\cdot|p_{2})$ and any watermark key $k\in\mathcal{K}$:
$\|F(P_{M}(\cdot|p_{1}),k)-F(P_{M}(\cdot|p_{2}),k)\|_{1}\leq L\cdot\|P_{M}(\cdot|p_{1})-P_{M}(\cdot|p_{2})\|_{1}.$ (7)
@@
Let $p_{1}$ and $p_{2}$ be two different prompts satisfying the similarity condition in Equation 5. Let $k_{1}$ and $k_{2}$ be two randomly sampled watermark keys from the key space $\mathcal{K}$. The effect of applying these keys on the output distribution should be highly consistent across prompts:
$\mathbb{E}_{k_{1},k_{2}\sim\mathcal{K}}[\text{Sim}(P_{M}^{F}(\cdot|p_{1},k_{1})-P_{M}^{F}(\cdot|p_{1},k_{2}),P_{M}^{F}(\cdot|p_{2},k_{1})-P_{M}^{F}(\cdot|p_{2},k_{2}))]\geq\rho$ (8)
@@
Based on Theorem 1, we calculate the average similarity using the estimated distributions from Step 2. To ensure stability across different sampling temperatures, we first apply a rank transformation. Specifically, for a token $y$, its rank is defined as the number of tokens with probability greater than or equal to that of $y$, denoted as $R(P)(y|x)=|\{y^{\prime}\in V:P(y^{\prime}|x)\geq P(y|x)\}|$. We then compute the expected similarity:
$\bar{S}=\frac{1}{N}\sum_{P_{i}\neq P_{j}\in\mathcal{P}}\sum_{k_{m}\neq k_{n}\in\mathcal{K}}\text{Sim}(\Delta_{R}(P_{i},k_{m},k_{n}),\Delta_{R}(P_{j},k_{m},k_{n}))$ (9)
Here, $\mathcal{P}$ is the prompt set, $\mathcal{K}$ is the watermark key set, $N=|\mathcal{P}|(|\mathcal{P}|-1)|\mathcal{K}|(|\mathcal{K}|-1)$, and $\Delta_{R}(P,k_{m},k_{n})=R(\hat{P}_{M}^{F}(\cdot|P,k_{m}))-R(\hat{P}_{M}^{F}(\cdot|P,k_{n}))$. We verify the importance of rank transformation in Appendix G.
@@ 4.3 Constructing Repeated Sampling With Same Watermark Key
Figure 2: Distribution of start keys for identical prefixes in KTH watermarking. Analysis based on prompts described in Section 4.3 for Watermark-Probe-v2. Each subplot represents a specific prefix(in title).
@@ 5.1 Experiment Setup
Water-Probe-v1 (w. prompt 4.3)
@@
Water-Probe-v2 (w. prompt 4.3)
@@ 6 Enhancing the Imperceptibility of Watermarked LLMs
Table 2 demonstrates that Water-Probe-v1 fails to effectively identify both Water-Bag and Exp-Edit algorithms. For Water-Probe-v2, detection difficulty increases with larger bag sizes in Water-Bag and longer key lengths in Exp-Edit. However, Water-Bag proves more challenging to identify. Crucially, Water-Bag’s detectability remains stable as $n$ increases, while Exp-Edit becomes more easy to identify. We analyzed the key distribution of Water-Bag under Water-Probe-v2 for different prefixes in Figure 4. The distribution is notably more uniform compared to Exp-Edit in Figure 2, explaining why Water-Bag is relatively harder to identify. We observe that increasing watermark key randomness reduces watermark robustness for both strategies, with Exp-Edit also significantly increasing detection time. This highlights a trade-off in watermarking algorithms between key randomness and robustness (Liu et al., 2024c). Future work should focus on developing algorithms that enhance randomness without compromising robustness.

### v3

@@ Abstract
Text watermarking for Large Language Models (LLMs) has made significant progress in detecting LLM outputs and preventing misuse. Current watermarking techniques offer high detectability, minimal impact on text quality, and robustness to text editing. However, current researches lack investigation into the imperceptibility of watermarking techniques in LLM services. This is crucial as LLM providers may not want to disclose the presence of watermarks in real-world scenarios, as it could reduce user willingness to use the service and make watermarks more vulnerable to attacks. This work investigates the imperceptibility of watermarked LLMs. We design the first unified identification method called Water-Probe that identifies all kinds of watermarking in LLMs through well-designed prompts. Our key motivation is that current watermarked LLMs expose consistent biases under the same watermark key, resulting in similar differences across prompts under different watermark keys. Experiments show that almost all mainstream watermarking algorithms are easily identified with our well-designed prompts, while Water-Probe demonstrates a minimal false positive rate for non-watermarked LLMs. Finally, we propose that the key to enhancing the imperceptibility of watermarked LLMs is to increase the randomness of watermark key selection. Based on this, we introduce the Water-Bag strategy, which significantly improves watermark imperceptibility by merging multiple watermark keys.
@@ 1 Introduction
Text watermarking techniques for LLMs have become an important method to mitigate the above issues by adding detectable features to LLM outputs (Liu et al., 2024b). Recent researches on LLM watermarking have focused on improving watermark detectability (Kirchenbauer et al., 2023a), minimizing impact on generated text (Aaronson & Kirchner, 2022), and enhancing robustness against text modifications (Liu et al., 2024a). However, no work has considered the imperceptibility of watermarked LLMs, i.e., whether users can know if an LLM service is watermarked. In real-world scenarios, LLM service providers may not disclose the existence of watermarks, as it could reduce user willingness to use the service and make it more vulnerable to attacks (Sadasivan et al., 2023). As more LLM services consider implementing watermarks, it is crucial to investigate whether users can identify watermarked LLMs solely through crafted prompts.
Some studies focus on the imperceptibility of watermarked text, ensuring watermarked and non-watermarked texts are indistinguishable (Hu et al., 2023; Wu et al., 2023b). However, even if individual watermarked texts are imperceptible, the distribution of numerous watermarked texts may reveal whether the LLM is watermarked, especially when repeatedly sampling with the same watermark key (Wu et al., 2024). While some studies explore cracking watermarks using large volumes of watermarked text (Jovanović et al., 2024; Sadasivan et al., 2023; Wu & Chandrasekaran, 2024), they assume the LLM is watermarked and cannot determine if the LLM is watermarked. The most relevant work is Gloaguen et al. (2024), which proposes a black-box detection method for watermarked LLMs. However, their approach uses different detection methods for different watermarks and cannot effectively detect all watermarking techniques.
@@
In this work, we propose Water-Probe, the first unified method for identifying watermarked LLMs that can detect all types of watermarks embedded during the LLM’s text generation process. (see related work section for this type of watermarking) Our motivation stems from a key observation: all current LLM watermarking algorithms expose consistent bias when repeatedly sampled under the same watermark key. Based on this, our Water-Probe algorithm first crafts prompts to perform repeated sampling under the same watermark key, then compares the consistency of sampling distribution differences across different prompts under a pair of watermark keys. Highly consistent differences indicate a watermarked LLM.
@@
Finally, we explore methods to enhance the imperceptibility of watermarked LLMs. We find that increasing the randomness of watermark key selection is crucial, as it makes it more difficult to construct prompts for repeated sampling using the same key. Based on this, we propose the Water-Bag algorithm, which combines multiple watermark keys into one, randomly selecting a key for each generation and choosing the highest score for detection. Although increasing key selection randomness often leads to a slight decrease in detectability, it significantly enhances the imperceptibility of watermarked LLMs. Addressing this trade-off should be an important direction for future work.
Our main contributions are summarized as follows:
  • We propose Water-Probe, the first unified algorithm that can detect various types of watermarked LLMs by analyzing the consistency of sampling distribution differences across different prompts under fixed watermark keys.
  • Through extensive experiments, we demonstrate that Water-Probe achieves high detection accuracy across different LLMs, watermarking methods, and sampling settings, while maintaining a low false positive rate for non-watermarked LLMs.
  • We introduce Water-Bag, a novel algorithm that enhances LLM watermark imperceptibility by combining multiple watermark keys, and analyze the trade-off between watermark detectability and imperceptibility.
@@ 2 Preliminaries
N-gram based watermarking ensures that for the same n-token prefix, the watermark key for generating the next token remains consistent. This approach is widely used in current watermarking algorithms, including KGW (Kirchenbauer et al., 2023a), KGW-V2 (Kirchenbauer et al., 2023b), Aar (Aaronson & Kirchner, 2022), DiPmark (Wu et al., 2023b), and SIR (Liu et al., 2024a). Next, we define fixed-key-list-based watermarking:
@@ 3.2 Pipeline of Watermarked LLM Identification
Observation 1 implies that the key to identifying a watermarked LLM is to construct prompts that allow for multiple samplings using the same watermark key to reveal the difference between $P_{M}^{F}(y_{i}|x,y_{1:i-1},k)$ and $P_{M}(y_{i}|x,y_{1:i-1})$. However, due to the black-box setting, we cannot directly access the origin logits $P_{M}(y_{i}|x,y_{1:i-1})$. Instead, we calculate the difference in LLM outputs for two distinct keys, defined as $\Delta(x,k_{m},k_{n})=P_{M}^{F}(\cdot|x,k_{m})-P_{M}^{F}(\cdot|x,k_{n})$, where we use $P_{M}^{F}(\cdot|x,k)$ to represent the output distribution.
If two prompts yield similar output distributions, the same watermark key should have similar effects on both prompts (proven in Theorem 1). We determine if an LLM contains a watermark by comparing the consistency of the effects of two watermark keys on two similar prompts. Specifically, given $x_{1}$, $x_{2}$, $k_{1}$, and $k_{2}$, we assess the similarity between $\Delta(x_{1},k_{1},k_{2})$ and $\Delta(x_{2},k_{1},k_{2})$. High similarity indicates the presence of a watermark; otherwise, we conclude there is no watermark. Based on the above analysis, we now present the process of the Water-Probe algorithm.
Step 1: Construct highly correlated prompts. Construct $N$ prompts $x_{1},x_{2},...,x_{N}$ such that their output probability distributions under $M$ are highly similar, which can be expressed as:
$\forall i,j\in\{1,2,...,N\},\text{KL}(P_{M}(\cdot|x_{i})||P_{M}(\cdot|x_{j}))\leq\epsilon\text{ and }x_{i}\neq x_{j}$ (5)
@@
Step 2: Sampling with simulated fixed watermark keys. Since we cannot access the logits under a given watermark key, we need to use repeated sampling to estimate the distribution. We construct a set of simulated watermark keys $K=\{k_{1},k_{2},...,k_{m}\}$ based on our prompt design (detailed in subsequent sections). For each prompt $x_{i}$ and each simulated key $k_{j}\in K$, we estimate the distribution as follows:
$\hat{P}_{M}^{F}(y|x_{i},k_{j})=\frac{1}{W}\sum_{w=1}^{W}\mathbf{1}_{y_{i,j}^{w}=y},\quad\text{where }y_{i,j}^{w}\sim P_{M}^{F}(y|x_{i},k_{j})$ (6)
where $W$ is the sample count, $\mathbf{1}_{A}$ is the indicator function, and $y_{i,j}^{w}$ is the $w$-th sample sampled from $P_{M}^{F}(y|x_{i},k_{j})$. Specific prompt techniques for different watermarking algorithms will be detailed later. Note that our prompt design for simulating watermark keys assumes the target LLM has a watermark. If it doesn’t, then $P_{M}^{F}(y|x_{i},k_{j})=P_{M}(y|x_{i})$ for all simulated keys.
@@
For prompts $x_{1}$ and $x_{2}$ satisfying the similarity condition in Equation 5, the watermark rule $F$ satisfies Lipschitz continuity. That is, there exists a constant $L>0$ such that for the probability distributions $P_{M}(\cdot|x_{1}),P_{M}(\cdot|x_{2})$ and any watermark key $k\in\mathcal{K}$:
$\|F(P_{M}(\cdot|x_{1}),k)-F(P_{M}(\cdot|x_{2}),k)\|_{1}\leq L\cdot\|P_{M}(\cdot|x_{1})-P_{M}(\cdot|x_{2})\|_{1}.$ (7)
@@
Let $x_{1}$ and $x_{2}$ be two different prompts satisfying the similarity condition in Equation 5. Let $k_{1}$ and $k_{2}$ be two randomly sampled watermark keys from the key space $\mathcal{K}$. The effect of applying these keys on the output distribution should be highly consistent across prompts:
$\mathbb{E}_{k_{1},k_{2}\sim\mathcal{K}}[\text{Sim}(P_{M}^{F}(\cdot|x_{1},k_{1})-P_{M}^{F}(\cdot|x_{1},k_{2}),P_{M}^{F}(\cdot|x_{2},k_{1})-P_{M}^{F}(\cdot|x_{2},k_{2}))]\geq\rho$ (8)
@@
Based on Theorem 1, we calculate the average similarity using the estimated distributions from Step 2. To ensure stability across different sampling temperatures, we first apply a rank transformation. Specifically, for a token $y$, its rank is defined as the number of tokens with probability greater than or equal to that of $y$, denoted as $R(P(y|x))=|\{y^{\prime}\in V:P(y^{\prime}|x)\geq P(y|x)\}|$. We then compute the expected similarity:
$\bar{S}=\frac{1}{N}\sum_{x_{i}\neq x_{j}\in\mathcal{X}}\sum_{k_{m}\neq k_{n}\in\mathcal{K}}\text{Sim}(\Delta_{R}(x_{i},k_{m},k_{n}),\Delta_{R}(x_{j},k_{m},k_{n}))$ (9)
Here, $\mathcal{X}$ is the prompt set, $\mathcal{K}$ is the watermark key set, $N=|\mathcal{X}|(|\mathcal{X}|-1)|\mathcal{K}|(|\mathcal{K}|-1)$, and $\Delta_{R}(x,k_{m},k_{n})=R(\hat{P}_{M}^{F}(\cdot|x,k_{m}))-R(\hat{P}_{M}^{F}(\cdot|x,k_{n}))$. We verify the importance of rank transformation in Appendix G.
@@ 3.3 Constructing Repeated Sampling With Same Watermark Key
Figure 2: Distribution of start keys for identical prefixes in Exp-Edit watermarking. Analysis based on prompts described in Section 3.3 for Watermark-Probe-v2. Each subplot represents a specific prefix(in title).
@@ 4.1 Experiment Setup
Water-Probe-v1 (w. prompt 3.3)
@@
Water-Probe-v2 (w. prompt 3.3)
@@ 5 Enhancing the Imperceptibility of Watermarked LLMs
Table 2 demonstrates that Water-Probe-v1 fails to effectively identify both Water-Bag and Exp-Edit algorithms. For Water-Probe-v2, detection difficulty increases with larger bag sizes in Water-Bag and longer key lengths in Exp-Edit. However, Water-Bag proves more challenging to identify. Crucially, Water-Bag’s detectability remains stable as $n$ increases, while Exp-Edit becomes more easy to identify. We analyzed the key distribution of Water-Bag under Water-Probe-v2 for different prefixes in Figure 4. The distribution is notably more uniform compared to Exp-Edit in Figure 2, explaining why Water-Bag is relatively harder to identify. We observe that increasing watermark key randomness reduces watermark robustness for both strategies, with Exp-Edit also significantly increasing detection time. This highlights a trade-off in watermarking algorithms between key randomness and robustness (Liu et al., 2024b). Future work should focus on developing algorithms that enhance randomness without compromising robustness.
6 Related Work
Large language model (LLM) watermarking techniques (Liu et al., 2024b) have become crucial for copyright protection (Sander et al., 2024), generated text detection (Wu et al., 2023a), and preventing misuse (Liu et al., 2024c). LLM watermarking can be broadly categorized into two types. The first type is post-processing watermarking, which modifies generated text using format-based (Sato et al., 2023), lexical-based, syntax-based (Wei et al., 2022), or generation-based (Zhang et al., 2024) approaches to add watermark. However, post-processing watermarking methods require waiting for text generation to complete before modifying and adding watermarks, which is not suitable for current LLM services that require real-time text generation. Another category of watermarking algorithms, known as watermarking during generation, typically involves adjusting the distribution of the next generated token based on a watermark key. For instance, the KGW (Kirchenbauer et al., 2023a) algorithm divides the vocabulary into red and green lists, increasing the probability of tokens in the green list. SIR (Liu et al., 2024a) further modifies logits based on semantic information, enhancing watermark robustness. An important objective of these methods is to maintain the imperceptibility of generated text, i.e., watermarked and non-watermarked text should have identical distributions, with some distortion-free algorithms showing promising results (Kuditipudi et al., 2023; Aaronson & Kirchner, 2022). However, previous work has overlooked the imperceptibility of watermarked LLMs themselves, i.e., whether external users can detect if an LLM service contains watermarks without disclosure. This work investigates the imperceptibility of watermarked LLMs. The most relevant work is (Gloaguen et al., 2024), our work differs in several aspects: we propose a unified identification method for all watermarking during generation algorithms without requiring different detection approaches for different watermarks. Our method also supports identification of complex watermark variants (like EXP-Edit with sampling), and we introduce the water-bag algorithm to enhance watermarked LLM imperceptibility.